# number of leading zeros ICWS improved consistent weighted sampling

### v1

@@ 1. Introduction
Distributed data processing and storage requires a strategy how tasks or data are mapped across available resources. In more abstract terms, the strategy must be able to distribute keys to a given number of buckets. Ideally, the keys are uniformly assigned to the buckets to achieve balanced utilization of resources. If there are $n$ buckets labeled $10,1,\ldots,n-$ 1, a very popular pattern to map a given integer key $k$ to a bucket index $b$ is
@@
If the key $k$ is the result of some high-quality hash function (Yi, [n.d.]; Vaneev, [n.d.]; Peters, [n.d.]) and can be considered like a true random value, and the number of resources is much smaller than the maximum possible hash value, which is usually the case for 64-bit hash values, the keys will be uniformly distributed over the buckets, such that
@@
Therefore, it is critical to minimize the number of reassignments, while preserving a uniform distribution. To maintain uniformity after increasing $n$ to $n+1$, an ideal reallocation would only move a minimal number of keys from the already existing $n$ buckets to the newly added $(n+1)$-th bucket having index $n$ to restore balance. The mapping of all other keys should remain untouched. The expected fraction of keys that have to be moved, when using an optimal balanced mapping strategy, must be at least $\frac{1}{n+1}$. Hence, the probability that a key gets reassigned would be close to 0, for large $n$. Algorithms that minimize the number of reassignments while mapping keys uniformly to a given number of buckets are known to be consistent. If there was no performance difference, it would be generally wise to replace pattern (1) by
$\operatorname*{ch}(k,n)\rightarrow b.$
where $\operatorname*{ch}(k,n)$ denotes a consistent hash function. This work presents JumpBackHash, which is a simple consistent hash function with a performance comparable to the modulo-based approach (1).
Table 1. Notation.
Symbol Comment
@@
$\mathcal{I}_{m}$ $\mathcal{I}_{m}:=\{2^{m},2^{m}+1,\ldots,2^{m+1}-1\}=\llbracket 2^{m+1}\rrbracket\setminus\llbracket 2^{m}\rrbracket$
@@ 1.1. Consistent Hash Functions
The probability that a random key is assigned to a bucket is the same for all $n$ buckets (Ioffe, 2010; Leu, 2023; Manasse et al., [n.d.])
@@
This property is sometimes also referred to as minimal disruption (Mendelson et al., 2021) or consistency (Ioffe, 2010). In this paper, we use the term consistency to describe the combination of monotonicity and uniformity.
While the modulo-based approach is uniform according to (2), it is not monotonic. A minimal example violating (3) is given by $k=4$ and $n=2$ when $\operatorname*{ch}(k,n):=k\bmod n$.
@@ 1.2. Related Work
Interestingly, consistent hashing can be seen as a special case of consistent weighted sampling (Manasse et al., [n.d.]), where the weight vector has a single dimension with an integer weight equal to the number of buckets $n$. The adaptation of the improved consistent weighted sampling (ICWS) algorithm (Ioffe, 2010) to our problem results in Algorithm 1 which obviously has a constant time complexity. Unfortunately, it involves expensive floating-point operations including an exponentiation and a logarithm evaluation, which makes it rather slow, especially on embedded or low-power devices without floating-point units. Nevertheless, the algorithm can be easily implemented, as it relies on an ordinary pseudorandom generator (PRG) that is seeded with the given key $k$. Standard libraries often already come with modern and reliable PRG implementations, such as SplitMix (Steele et al., [n.d.]) in Java, that pass a series of statistical tests (Lemire, [n.d.]). These libraries sometimes even offer functions for the efficient generation of exponentially distributed random values, as required by ICWS. Although ICWS has been around for some time, it was only recently discovered as a consistent hash function when it was added to version 0.14.0 of our open source Hash4j library (Has, [n.d.]).
@@
JumpHash (Lamping and Veach, 2014), on the other hand, became much more popular, although it was invented a few years later. It is shown in Algorithm 2 and will be discussed in more detail later. As ICWS, it uses a standard PRG and floating-point arithmetic, but avoids expensive transcendental functions. Therefore, it is usually faster than ICWS, except for large $n$ due to its logarithmic time complexity. JumpHash is widely used in the meantime. For example, it is offered as hash function in the Guava Java library (Gua, [n.d.]) or in ClickHouse (Cli, [n.d.]). It is also used by Booking.com for their customer review system (Hiltpolt, 2022), by Grafana for trace and metric data distribution (Agarwal, 2020; Boucault, 2020), by VMware Greenplum since version 6 to map data values to segments (Gre, [n.d.]), by RabbitMQ since version 3.7.8 in its Consistent Hashing Exchange Plugin for message distribution (Rab, [n.d.]), or for distributed processing of vehicle license plates (Wang et al., 2023).
@@
Recently, PowerHash was introduced as the first consistent hash algorithm with constant runtime able to beat JumpHash over the whole range (Leu, 2023). Unlike JumpHash and ICWS, it relies on a family of hash functions that are less common in standard libraries and often have to be self-implemented. Since families of hash functions are used less frequently in practice, their statistical properties such as independence and uniformity are usually less tested and therefore less trustworthy than those of PRGs. A family of hash functions is also used by the even more recent FlipHash algorithm (Masson and Lee, 2024), which is similar to PowerHash but avoids floating-point operations.
1.3. Our Contributions
To derive JumpBackHash, we use the concept of active indices, which has already been used for consistent weighted sampling (Manasse et al., [n.d.]; Ioffe, 2010). This new perspective allows a better understanding of JumpHash and leads us after some stepwise optimizations to JumpBackHash. We also perform a theoretical runtime analysis, which is empirically verified together with uniformity and monotonicity.
@@ 1.3. Our Contributions
        $\triangleright$ $G$ is $\operatorname*{Gamma}(2,1)$-distributed
@@
Algorithm 1 Specialization of improved consistent weighted sampling (ICWS) (Ioffe, 2010) to map a key $k$ consistently to $\llbracket n\rrbracket$.
2. Methodology
A simple way to define a consistent hash function is
Unfortunately, this approach is not very efficient as it takes $\mathcal{O}(n)$ time to generate all relevant random values $R_{b}$ with $b\in\llbracket n\rrbracket$. The order in which the random values $R_{b}$ are drawn using the PRG is not important as long as it is well defined and does not depend on $n$. For example, if we know an upper bound $N$ such that $n\leq N$ is always satisfied, we could draw the random values in reverse order according to $R_{N-1},R_{N-2},\ldots,R_{1},R_{0}$. Even though this would lead to a time complexity of $\mathcal{O}(N)$ if $n\ll N$, (4) would still be a valid consistent hash function.
2.1. Active Indices
To construct a more efficient consistent hash algorithm based on (4), we use the concept of active indices (Manasse et al., [n.d.]; Ioffe, 2010). A bucket index $b$ is an active index if
(5) $R_{b}=\min_{b^{\prime}\in\llbracket b+1\rrbracket}R_{b^{\prime}}.$
@@ 2.1. Active Indices
Algorithm 2 The JumpHash algorithm (Lamping and Veach, 2014) consistently maps $k$ to the integer range $\llbracket n\rrbracket$.
@@ 2.2. JumpHash
Generating active indices in ascending order is exactly what JumpHash does, even though the algorithm was derived and analyzed without the concept of active indices (Lamping and Veach, 2014). Given an active index $A$, the probability that $A+1$ is the next larger active index equals the probability that $R_{A+1}$ is the smallest among the first $A+2$ random values $\{R_{b}\}_{b\in\llbracket A+2\rrbracket}$ which occurs with a probability of $\frac{1}{A+2}$. The probability that $A+2$ is the next active index with respect to $A$ is given by the probability that $A+1$ is not an active index multiplied by the probability that $R_{A+2}$ is the smallest value, which is therefore $(1-\frac{1}{A+2})\frac{1}{A+3}=\frac{A+1}{(A+2)(A+3)}$. In general, the probability that $A+\Delta$ is the next active index with respect to $A$ is given by $(\prod_{i=2}^{\Delta}1-\frac{1}{A+i})\frac{1}{A+\Delta+1}=\frac{A+1}{(A+\Delta)(A+\Delta+1)}$.
@@
$A_{l+1}=\lfloor(A_{l}+1)/U\rfloor\qquad\text{with $U\sim\operatorname*{Uniform}([0,1))$ and $A_{0}=0$,}$
@@
where $\operatorname*{Bernoulli}(p)$ denotes a Bernoulli random variable with probability $p$. When using $\operatorname*{\mathbb{E}}(X+Y)=\operatorname*{\mathbb{E}}(X)+\operatorname*{\mathbb{E}}(Y)$ and $\operatorname*{Var}(X+Y)=\operatorname*{Var}(X)+\operatorname*{Var}(Y)$ for independent $X$ and $Y$ together with $\operatorname*{\mathbb{E}}(\operatorname*{Bernoulli}(p))=p$ and $\operatorname*{Var}(\operatorname*{Bernoulli}(p))=p(1-p)$, the expectation and the variance of $T_{\text{JBH}}$ can be derived as
$\operatorname*{\mathbb{E}}$ $(T_{\text{JH}})$ $=\sum_{i=1}^{n}\frac{1}{i}$ $\leq 1+\ln n,$
$\operatorname*{Var}$ $(T_{\text{JH}})$ $=\sum_{i=1}^{n}\frac{1}{i}-\sum_{i=1}^{n}\frac{1}{i^{2}}$ $\leq\ln n.$
@@ 2.3. Jumping Backwards
The way JumpHash generates active indices is not unique. If an upper bound $N$ is known for the number of buckets $n\leq N$, which is always the case in practice as standard integer types are limited, it is also possible to generate active indices in descending order using the recursion
$A_{l+1}\sim\operatorname*{Uniform}(\llbracket A_{l}\rrbracket)\qquad\text{with$A_{0}=N$.}$
The rationale behind this approach is very simple. The next smallest active index is given by the smallest random value from $\{R_{b}\}_{b\in\llbracket A_{l}\rrbracket}$. As, by definition, they are all identically distributed, they have the same chance to be the next smaller active index. Based on this recursion another consistent hash algorithm can be constructed as shown by Algorithm 3.
@@
In contrast to JumpHash, this approach does not involve floating-point operations. However, it requires the sampling from an integer range (O’Neil, 2018; Lemire, 2019). Another downside is that all active indices in $[n,N)$ must be generated. If $n\ll N$, which is usually the case, the runtime complexity would be $\mathcal{O}(\log N)$, which is worse than the $\mathcal{O}(\log n)$ of JumpHash. While we prefer the absence of floating-point operations, we need to make the generation of active indices in descending order much more efficient.
@@ 2.4. Improved Active Index Generation
A good reason to consider intervals with power-of-two cardinalities is that uniform integer values can be efficiently sampled from $\mathcal{I}_{m}$ using $m$ random bits. The intervals $\mathcal{I}_{m}$ can be ordered as all values of $\mathcal{I}_{m}$ are smaller than that of $\mathcal{I}_{m+1}$ and also the cardinality of $\mathcal{I}_{m}$ is smaller than that of $\mathcal{I}_{m+1}$. If we compare those intervals in the following, we always refer to this ordering.
To obtain active indices in descending order, we process the intervals $\mathcal{I}_{m}$ with $m\in\llbracket M\rrbracket$ in descending order and generate all active indices therein also in descending order. The last active index is always 0 by definition (5). The probability that $\mathcal{I}_{m}$ contains an active index is 50% as the next smaller random value belongs either to $\{R_{b}\}_{b\in\llbracket 2^{m}\rrbracket}$ or $\{R_{b}\}_{b\in\mathcal{I}_{m}}$ with the same probability. This decision requires a single random bit $X_{m}\in\{0,1\}$. In 50% of all cases we are done with $\mathcal{I}_{m}$, otherwise we actually have to generate the occurring active indices within $\mathcal{I}_{m}$.
Since we know at this point that there is at least one active index in $\mathcal{I}_{m}$, the largest active index can simply be generated by drawing a uniform random value $Y_{m}$ from $\mathcal{I}_{m}$. It is possible that $\mathcal{I}_{m}$ contains further active indices in the interval. As the next smaller active index is uniformly distributed over $\llbracket Y_{m}\rrbracket$, we sample a uniform random value $Z_{m,0}$ from $\llbracket Y_{m}\rrbracket$. If $Z_{m,0}\notin\mathcal{I}_{m}$, $Z_{m,0}$ is rejected as active index and the process is continued with the next smaller interval $\mathcal{I}_{m-1}$. Otherwise, $Z_{m,0}$ is the next smaller active index within $\mathcal{I}_{m}$ and we generate further random values according to
$Z_{m,v}\sim\operatorname*{Uniform}(\llbracket Z_{m,v-1}\rrbracket),$
@@
The whole procedure to generate active indices in descending order, which is statistically equivalent to Algorithm 3, is summarized in Algorithm 4. While $X_{m}$ and $Y_{m}$ are already very convenient to sample, $Z_{m,v}$ needs to be sampled from a non-power-of-two interval $\llbracket b\rrbracket$. One possibility would be to use Lemire’s method (Lemire, 2019). A simpler and also unbiased approach that fits better in our case is to sample from the smallest enclosing power-of-two interval, which is by definition $\llbracket 2^{m+1}\rrbracket$, and reject values that are out of range (O’Neil, 2018).
Applying the rejection method instead of sampling from $\llbracket b\rrbracket$ is statistically equivalent to using option 2 in Algorithm 4. With that, all required random values are either booleans or from a power-of-two interval and can therefore be generated efficiently.
@@
       for $m\leftarrow M-1$ to 00 do
@@
             $Y_{m}\leftarrow 2^{m}+\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{m}\rrbracket)]$
              $\triangleright$ sample $Y_{m}$ from $\mathcal{I}_{m}$
@@
                   $Z_{m,v}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket b\rrbracket)]$
                    $\triangleright$ option 1
@@
Algorithm 4 Consisent hashing based on a more efficient generation of active indices in reverse order, $n\leq 2^{M}$.

### v2

# improved consistent weighted sampling PRG pseudorandom generator

@@ Abstract.
consistent hashing, sharding, distributed storage, distributed processing
@@ 1. Introduction
Distributed data processing and storage requires a strategy how tasks or data are mapped across available resources. In more abstract terms, the strategy must be able to distribute keys to a given number of buckets. Ideally, the keys are uniformly assigned to the buckets to achieve balanced utilization of resources. If there are $n$ buckets labeled $10,1,\ldots,n-$ 1, a very popular pattern to map a given integer key $k$ to a bucket index $b$ uses the modulo operation according to
@@
If the key $k$ is the result of some high-quality hash function (Yi, [n. d.]; Vaneev, [n. d.]; Peters, [n. d.]) and can be considered like a true random value, and the number of buckets is much smaller than the maximum possible hash value, which is usually the case for 64-bit hash values, the keys will be uniformly distributed over the buckets, such that
@@
Therefore, it is critical to minimize the number of reassignments while preserving a uniform distribution. To maintain uniformity after increasing the number of buckets from $n$ to $n+1$, some keys must be shifted from the already existing $n$ buckets to the newly added $(n+1)$-th bucket with index $n$. Since the expected proportion of keys in the new bucket should be $\frac{1}{n+1}$ at the end, it is obvious that on average at least $\frac{1}{n+1}$ of the keys must switch buckets. Consistent hash algorithms achieve this lower bound and thus minimize the number of reassignments while mapping keys uniformly to a given number of buckets. For them, the probability that a key gets reassigned would be close to 0 for large $n$, because $\frac{1}{n+1}\rightarrow 0$ as $n\rightarrow\infty$. If there was no performance difference, it would be generally wise to replace pattern (1) by
$\operatorname*{ch}(k,n)\rightarrow b$
where $\operatorname*{ch}(k,n)$ denotes a consistent hash function.
Figure 1 illustrates the difference between the modulo-based approach and consistent hashing in terms of the number of reassignments when the number of buckets changes from 3 to 4. This work presents JumpBackHash, which is a simple consistent hash function with a performance comparable to the modulo-based approach (1). Table 1 provides an overview of the used notations.
Figure 1. When changing the number of buckets, some keys (letters) must be mapped to different buckets to restore balance. In contrast to the modulo operation, which puts almost all keys into different buckets, consistent hashing minimizes the expected number of reassignments (arrows).
Table 1. Notations.
Symbol Description
@@
$N$ some constant upper bound such that $n\leq N$ holds, see Section 2.3
$M$ some constant such that $n\leq 2^{M}$ holds, see Section 2.4
@@
$\sim$ is distributed as, e.g. $\operatorname*{Bernoulli}(0.5)\sim\operatorname*{Uniform}(\{0,1\})$
$\Pr$ probability, e.g. $\Pr(\operatorname*{Bernoulli}(0.3)=1)=0.3$
$\operatorname*{\mathbb{E}}$ mean, e.g. $\operatorname*{\mathbb{E}}(\operatorname*{Uniform}(\{0,1\}))=0.5$
$\operatorname*{Var}$ variance
@@
$\mathcal{I}_{m}$ $\mathcal{I}_{m}:=\{2^{m},2^{m}+1,\ldots,2^{m+1}-1\}=\llbracket 2^{m+1}\rrbracket\setminus\llbracket 2^{m}\rrbracket$, $m\geq 0$
$\mathcal{O}$ big O notation, e.g. $\sum_{i=1}^{n}\frac{1}{i}=\mathcal{O}(\ln n)$
$\alpha_{n}$ $\alpha_{n}:=\frac{1}{n}2^{\lfloor\log_{2}(n-1)\rfloor+1}\in[1,2)$, $n\geq 2$
@@ 1.1. Consistent Hash Functions
The probability that a random key is assigned to a bucket is the same for all $n$ buckets which can be written as (Ioffe, 2010; Leu, 2023; Manasse et al., 2010)
@@
This property is sometimes also referred to as minimal disruption (Mendelson et al., 2021) or consistency (Ioffe, 2010). In contrast, in this paper, we use the term consistency to describe the combination of monotonicity and uniformity.
While the modulo-based approach is uniform according to (2), it is not monotonic. A minimal example violating (3) is given by $k=4$ and $n=2$ when $\operatorname*{ch}(k,n):=k\bmod n$. The implication in (3) is actually an equivalence, because $\operatorname*{ch}(k,n+1)=n>\operatorname*{ch}(k,n)\in\llbracket n\rrbracket$. Therefore, the probability of reassignments for consistent hash functions is given by $\Pr(\operatorname*{ch}(k,n)\neq\operatorname*{ch}(k,n+1))=\Pr(\operatorname*{ch}(k,n+1)=n)=\frac{1}{n+1}$. Hence, uniformity and monotonicity are sufficient to minimize the expected number of reassignments.
@@ 1.2. Related Work
Interestingly, consistent hashing can be seen as a special case of consistent weighted sampling (Manasse et al., 2010), where the weight vector has a single dimension with an integer weight equal to the number of buckets $n$. The adaptation of the improved consistent weighted sampling (ICWS) algorithm (Ioffe, 2010) to our problem results in Algorithm 1 which obviously has a constant time complexity. Unfortunately, the algorithm involves expensive floating-point operations including an exponentiation and a logarithm evaluation, which makes it rather slow, especially on embedded or low-power devices without floating-point units. Nevertheless, the algorithm can be easily implemented, as it relies on an ordinary pseudorandom generator (PRG) that is seeded with the given key $k$. Standard libraries often already come with modern and reliable PRG implementations, such as SplitMix (Steele et al., 2014) in Java, that pass a series of statistical tests (Lemire, [n. d.]). These libraries sometimes even offer functions for the efficient generation of exponentially distributed random values, as required by ICWS. Although ICWS has been around for some time, it was only recently discovered as a consistent hash function when it was added to version 0.14.0 of our open source Hash4j library (Has, [n. d.]).
@@
JumpHash (Lamping and Veach, 2014), on the other hand, became much more popular, although it was invented a few years afterwards. It is shown in Algorithm 2 and will be discussed in more detail later. As ICWS, it uses a standard PRG and floating-point arithmetic, but avoids expensive transcendental functions. Therefore, it is usually faster than ICWS, except for large $n$ due to its logarithmic time complexity. JumpHash is widely used in the meantime. For example, it is offered as hash function in the Guava Java library (Gua, [n. d.]) or in ClickHouse (Cli, [n. d.]). It is also used by Booking.com for their customer review system (Hiltpolt, 2022), by Grafana for trace and metric data distribution (Agarwal, 2020; Boucault, 2020), by VMware Greenplum since version 6 to map data values to segments (Gre, [n. d.]), by RabbitMQ since version 3.7.8 in its Consistent Hashing Exchange Plugin for message distribution (Rab, [n. d.]), or for distributed processing of vehicle license plates (Wang et al., 2023).
@@
Recently, PowerHash was introduced as the first consistent hash algorithm with constant runtime able to beat JumpHash over the whole range of bucket numbers (Leu, 2023). Unlike JumpHash and ICWS, it relies on a family of hash functions that are less common in standard libraries and often have to be self-implemented. Since families of hash functions are used less frequently in practice, their statistical properties such as independence and uniformity are usually less tested and therefore less trustworthy than those of PRGs. A family of hash functions is also used by the even more recent FlipHash algorithm (Masson and Lee, 2024), which is similar to PowerHash but avoids floating-point operations.
@@
        $\triangleright$ $G$ follows a $\operatorname*{Gamma}(2,1)$ distribution
@@
Algorithm 1 Specialization of improved consistent weighted sampling (ICWS) (Ioffe, 2010) to map a key $k$ consistently to $\llbracket n\rrbracket:=\{0,1,2,\ldots,n-1\}$.
@@
              $\triangleright$ determine next larger active index, compare (7)
@@
Algorithm 2 The JumpHash algorithm (Lamping and Veach, 2014) consistently maps $k$ to the integer range $\llbracket n\rrbracket:=\{0,1,2,\ldots,n-1\}$.
1.3. Our Contributions
To derive JumpBackHash, we use the concept of active indices, which has already been used for consistent weighted sampling (Manasse et al., 2010; Ioffe, 2010). This new perspective allows a better understanding of JumpHash and leads us after some stepwise optimizations to JumpBackHash. We also perform a theoretical runtime analysis, which is empirically verified together with uniformity and monotonicity.
2. Methodology
A simple way to define a consistent hash function is
Unfortunately, this approach is not very efficient as it takes $\mathcal{O}(n)$ time to generate all relevant random values $R_{b}$ with $b\in\llbracket n\rrbracket$. The order in which the random values $R_{b}$ are drawn using the PRG is not important as long as it is well defined and does not depend on $n$. For example, if we know an upper bound $N$ such that $n\leq N$ is always satisfied, we could draw the random values in reverse order according to $R_{N-1},R_{N-2},\ldots,R_{1},R_{0}$. Even though this would lead to a much worse time complexity of $\mathcal{O}(N)$ if $n\ll N$, (4) would still be a valid consistent hash function.
2.1. Active Indices
To construct a more efficient consistent hash algorithm based on (4), we use the concept of active indices (Manasse et al., 2010; Ioffe, 2010). A bucket index $b$ is an active index if it satisfies
(5) $R_{b}=\min_{b^{\prime}\in\llbracket b+1\rrbracket}R_{b^{\prime}}.$
@@ 2.2. JumpHash
Generating active indices in ascending order is exactly what JumpHash does, even though the algorithm was derived and analyzed without the concept of active indices (Lamping and Veach, 2014). Given an active index $A$, the probability that $A+1$ is the next larger active index equals the probability that $R_{A+1}$ is the smallest among the first $A+2$ random values $\{R_{b}\}_{b\in\llbracket A+2\rrbracket}$ which occurs with a probability of $\frac{1}{A+2}$. The probability that $A+2$ is the next active index relative to $A$ is given by the probability that $A+1$ is not an active index multiplied by the probability that $R_{A+2}$ is the smallest value, which is therefore $(1-\frac{1}{A+2})\frac{1}{A+3}=\frac{A+1}{(A+2)(A+3)}$. In general, the probability that $A+\Delta$ is the next active index relative to $A$ is given by $(\prod_{i=2}^{\Delta}1-\frac{1}{A+i})\frac{1}{A+\Delta+1}=\frac{A+1}{(A+\Delta)(A+\Delta+1)}$.
@@
(7) $A_{l+1}=\lfloor(A_{l}+1)/U\rfloor\qquad\text{with $U\sim\operatorname*{Uniform}([0,1))$ and $A_{0}=0$,}$
@@
where $\operatorname*{Bernoulli}(p)$ denotes a Bernoulli random variable with probability $p$. When using
$\operatorname*{\mathbb{E}}\!\left(\sum_{i=1}^{n}K_{i}\right)$ $=\sum_{i=1}^{n}\operatorname*{\mathbb{E}}(K_{i}),$
$\operatorname*{Var}\!\left(\sum_{i=1}^{n}K_{i}\right)$ $=\sum_{i=1}^{n}\operatorname*{Var}(K_{i})$
for independent $K_{i}$ together with the known expressions for Bernoulli distributed random variables
(8) $\operatorname*{\mathbb{E}}(\operatorname*{Bernoulli}(p))$ $=p,$
(9) $\operatorname*{Var}(\operatorname*{Bernoulli}(p))$ $=p(1-p),$
the mean and the variance of $T_{\text{JBH}}$ can be derived as
(10) $\operatorname*{\mathbb{E}}$ $(T_{\text{JH}})$ $=\sum_{i=1}^{n}\frac{1}{i}$ $\leq 1+\ln n,$
(11) $\operatorname*{Var}$ $(T_{\text{JH}})$ $=\sum_{i=1}^{n}\frac{1}{i}\left(1-\frac{1}{i}\right)=\left(\sum_{i=1}^{n}\frac{1}{i}\right)-\left(\sum_{i=1}^{n}\frac{1}{i^{2}}\right)$ $\leq\ln n.$
This also shows the logarithmic runtime complexity $\mathcal{O}(\ln n)$ of JumpHash.
@@ 2.3. Jumping Backwards
The way JumpHash generates active indices (7) is not unique. If an upper bound $N$ is known for the number of buckets $n\leq N$, which is always the case in practice as standard integer types are limited, it is also possible to generate active indices in descending order using the recursion
(12) $A_{l+1}\sim\operatorname*{Uniform}(\llbracket A_{l}\rrbracket)\qquad\text{with$l\geq 0$ and $A_{0}=N$.}$
The rationale behind this approach is very simple. The next smallest active index is given by the smallest random value from $\{R_{b}\}_{b\in\llbracket A_{l}\rrbracket}$. As, by definition, the values in this set are all identically distributed, they all have the same chance of being the smallest. As a result, the next smaller active index is uniformly distributed over $\llbracket A_{l}\rrbracket$. Based on this recursion (12) another consistent hash algorithm can be constructed as shown by Algorithm 3.
@@
              $\triangleright$ determine next smaller active index, compare (12)
@@
In contrast to JumpHash, this approach does not involve floating-point operations. However, it requires sampling from an integer range (O’Neil, 2018; Lemire, 2019). Another downside is that all active indices in $[n,N)$ must be generated. If $n\ll N$, which is usually the case, the runtime complexity would be $\mathcal{O}(\log N)$, which is worse than the $\mathcal{O}(\log n)$ of JumpHash. While we prefer the absence of floating-point operations, we need to make the generation of active indices in descending order much more efficient.
@@ 2.4. Improved Active Index Generation
A good reason to consider intervals with power-of-two cardinalities is that uniform integer values can be efficiently sampled from $\mathcal{I}_{m}$ using exactly $m$ random bits. The intervals $\mathcal{I}_{m}$ can be ordered as all values of $\mathcal{I}_{m}$ are smaller than that of $\mathcal{I}_{m+1}$ and also the cardinality of $\mathcal{I}_{m}$ is smaller than that of $\mathcal{I}_{m+1}$. If we compare those intervals in the following, we always refer to this ordering. In particular, this allows us to say that $\mathcal{I}_{m}$ is smaller than $\mathcal{I}_{m^{\prime}}$ if $m<m^{\prime}$.
To obtain active indices in descending order, we process the intervals $\mathcal{I}_{m}$ with $m\in\llbracket M\rrbracket$ in descending order and generate all active indices therein also in descending order. The last active index is always 0 by definition (5). The probability that $\mathcal{I}_{m}$ contains an active index is 50% as the next smaller random value belongs either to $\{R_{b}\}_{b\in\llbracket 2^{m}\rrbracket}$ or $\{R_{b}\}_{b\in\mathcal{I}_{m}}$ with the same probability. This decision requires a single random bit $X_{m}\in\{0,1\}$. If $X_{m}=0$, which occurs in 50% of all cases, we know that there is no active index in $\mathcal{I}_{m}$ and we can continue processing the next smaller interval. Otherwise, if $X_{m}=1$, there is at least one active index, but we still need to find out which indices from $\mathcal{I}_{m}$ are actually active.
As all indices $b\in\mathcal{I}_{m}$ have the same chance that their associated random value $R_{b}$ is the smallest among them, the largest active index in $\mathcal{I}_{m}$ can simply be generated by drawing a uniform random value $Y_{m}$ from $\mathcal{I}_{m}$. It is possible that $\mathcal{I}_{m}$ contains further (smaller) active indices in the interval. As the next smaller active index (not necessarily in $\mathcal{I}_{m}$) is uniformly distributed over $\llbracket Y_{m}\rrbracket$, we sample a uniform random value $Z_{m,0}$ from $\llbracket Y_{m}\rrbracket$ analogous to (12). If $Z_{m,0}\notin\mathcal{I}_{m}$, $Z_{m,0}$ is rejected as active index and the process is continued with the next smaller interval $\mathcal{I}_{m-1}$. Otherwise, $Z_{m,0}$ is the next smaller active index within $\mathcal{I}_{m}$, and we iteratively generate further random values according to
(13) $Z_{m,v}\sim\operatorname*{Uniform}(\llbracket Z_{m,v-1}\rrbracket)\qquad\text{with $v\geq 1$},$
@@
The whole procedure to generate active indices in descending order, which is statistically equivalent to Algorithm 3, is summarized in Algorithm 4. While $X_{m}$ and $Y_{m}$ are already very convenient to sample, $Z_{m,v}$ needs to be sampled from a non-power-of-two interval $\llbracket b\rrbracket$. One possibility would be to use Lemire’s method (Lemire, 2019). A simpler and also unbiased approach that fits better in our case is to sample from the smallest enclosing power-of-two interval, which is by definition $\llbracket 2^{m+1}\rrbracket$ for $\mathcal{I}_{m}$, according to
(14) $Z_{m,v}\sim\operatorname*{Uniform}(\llbracket 2^{m+1}\rrbracket)\qquad\text{with $v\geq 1$}.$
As the sequence $(Z_{m,v})_{v\geq 0}$ is no longer strictly monotonically decreasing, we must ignore values that are not smaller than all previous values to ensure equivalent behavior, when replacing (13) in the inner loop of Algorithm 4, labeled as option 1, by (14), labeled as option 2. Fortunately, the inner loop takes care of this without any further modifications. If there is a value that fulfills one of the two stop conditions, but there was already another smaller or equal value before, this other value would have stopped the loop earlier. We continue with option 2, as all required random values are then either Boolean values or values from a power-of-two interval and can therefore be generated easily and efficiently by taking a corresponding number of random bits.
@@
       for $m\leftarrow M-1$ to 00 do $\triangleright$ process intervals $\mathcal{I}_{M-1},\mathcal{I}_{M-2},\ldots,\mathcal{I}_{0}$ in descending order
@@
              $\triangleright$ continue with next interval if $\mathcal{I}_{m}$ does not contain any active indices $Y_{m}\leftarrow 2^{m}+\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{m}\rrbracket)]$
              $\triangleright$ sample largest active index $Y_{m}$ from $\mathcal{I}_{m}$
@@
                    $\triangleright$ the first active index in $\mathcal{I}_{m}$ smaller than $n$ is returned $Z_{m,v}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket b\rrbracket)]$
                    $\triangleright$ option 1, compare (13)
@@
                    $\triangleright$ stop if there are no further active indices in $\mathcal{I}_{m}$
@@
Algorithm 4 Consistent hashing based on a more efficient generation of active indices in reverse order, $n\leq 2^{M}$.